OpenStudy (anonymous):

What is the measure of the angle formed by the minute hand and the hour hand of a clock when the clock shows 4:35? A.72.5 degrees B.90 degrees C. 120 degrees D. 72 degrees

OpenStudy (anonymous):

the answer is B?. i assume u mean angle in between hour hand and minute hand. angle for minute hand = 35/60 * 360 = 210 Hour hand = 4/12 * 360= 120. angle diff= 210-120= 90 degrees

OpenStudy (anonymous):

answer isnt B .... B is wrong

OpenStudy (anonymous):

The Right Answer Would Be A